# Deformable Convolution Module with Globally Learned Relative Offsets for Fundus Vessel Segmentation

### v1

@@ Abstract
Deformable convolution can adaptively change the shape of convolution kernel by learning offsets to deal with complex shape features. We propose a novel plug-and-play deformable convolutional module that uses attention and feedforward networks to learn offsets, so that the deformable patterns can capture long-distance global features. Compared with previously existing deformable convolutions, the proposed module learns the sub-pixel displacement field and adaptively warps the feature maps across all channels (rather than directly deforms the convolution kernel), which is equivalent to a relative deformation of the kernel’s sampling grids, achieving global feature deformation and the decoupling of (kernel size)–(learning network). Considering that the fundus blood vessels have globally self-similar complex edges, we design a deep learning model for fundus blood vessel segmentation, GDCUnet, based on the proposed convolutional module. Empirical evaluations under the same configuration and unified framework show that GDCUnet has achieved state-of-the-art performance on public datasets. Further ablation experiments demonstrated that the proposed deformable convolutional module could more significantly learn the complex features of fundus blood vessels, enhancing the model’s representation and generalization capabilities. The proposed module is similar to the interface of conventional convolution, we suggest applying it to more machine vision tasks with complex global self-similar features. Our source code will be available at https://github.com/LexyZhu/GDCUnet.git.
@@ 3.1 SAFDConvolution
    $\mathbf{K_{s}}=3,\;\mathbf{D_{s}}=1\;\Rightarrow\;$ $3\times 3$ grid $\{(-1,-1),(-1,0),(0,-1),\dots,(1,1)\}$;
@@
    $\mathbf{K_{s}}=3,\;\mathbf{D_{s}}=2\;\Rightarrow\;$ $5\times 5$ grid $\{(-2,-2),(-2,0),(0,-2),\dots,(2,2)\}$.
@@
The number of channels $C$ is usually divisible by $2$, we denote that $M_{c}=C/2$ and split the last dimension:
@@
$\Delta\mathbf{p}(\mathbf{p}_{0}+\mathbf{p}_{n})$ $=\frac{1}{M_{c}}\sum_{m_{c}=1}^{M_{c}}\tilde{X}_{offset}$ $\in\mathbb{R}^{2};$ (21)
@@ 3.2 GDCUnet
During the encoding stage, the Conv Block contains two $3\times 3$ conventional convolutions. SAFDConv Block uses two consecutive proposed SAFDConvolutions, where the hyperparameters can be set according to Table 1. Further, the Feature Incentive Block uses one $7\times 7$ conventional convolution.
@@
Figure 3: Overall pipeline of GDCUnet, where $C_{s}$ is $16$. Our network adopts a structure similar to Unet. In the encoding stage, Max-pooling is used to reduce the size of the feature map, and in the decoding stage, bilinear interpolation is used to restore the size of the feature map. We use tensor addition to achieve feature fusion of the feature maps from the previous stage. The modules using proposed SAFDConvolution are italicized bold.. To the best of our knowledge, the proposed deformable convolution is more suitable in the intermediate stage of feature extraction and decoding.
@@ 4.2 Implementation Details
Our experiments are built on CHASEDB1 [36], a public retinal blood vessel reference dataset. The dataset is divided into a training set and a test set in a ratio of $0.86:0.14$. The experimental configuration summary is presented in Table 2. All experiments are conducted using the Pytorch framework [37], and the learning rate is uniformly optimized using cosine annealing. We evaluate the model on the test set after each epoch and report the highest score achieved.
@@
Input Resolution $256\times 256$
Output Resolution $256\times 256$
@@ 4.3 Metrics
Table 3: Summary of Evaluation Metrics.
@@ 4.4 Main Results
Rolling-Unet #1.78M [9] AAAI’2024 0.6008 0.7506 15.00 1.414 0.7745 0.9828 0.7568
@@
Knowledge  [^2] We emphasize conducting experiments under the unified benchmark and configuration. However, the model based on prior knowledge characterizing the vascular topological structure involves separate settings for data processing, loss functions, etc., and experiments need to be conducted under separate configurations. DSCNet #- -M [34] ICCV’2023 0.6073 0.7521 15.00 2.0000 0.7534 0.9832 0.7638
@@
The results of all models indicate that under the same configuration, this is a challenging project. Almost all GDCUNet Settings have achieved competitive performance. Setting 5 achieves the best results in multiple metrics such as IoU and Dice, with only 1.40M parameters. At the same time, there is also more lightweight option (Setting 3) that achieves a balance between parameter quantity and segmentation performance. DconnNet has the highest recall but is limited overall. Benefit by the characterization ability of SAFDConvolution for globally self-similar complex local features, GDCUNet Setting 5 has reached state-of-the-art comprehensive level.
@@ 5.1 The Effect of SAFDConvolution
The main improvement made by GDCUNet is the addition of SAFDConvolution during the feature processing, thereby more effectively capturing complex local details with global self-similarity. Compared with other existing deformable convolution methods, SAFDConvolution can handle global features more effectively. We replace the convolution in the SAFDConvolution Block with other convolution (unify the kernel size $5\times 5$), and the quantitative evaluation results are shown in Table 5.
@@ 5.3 T-SNE Visualization
To further verify the generalization performance of SAFDConvolution, following the approach of J. Feng et al. [43], we use t-SNE to reduce the dimension of learned feature tensors. First, visualize the domain-invariant features extracted by SAFDConvolutions in GDCUnet. Then, replace them with conventional convolutions while keeping other structures unchanged to visualize the domain-invariant features extracted by conventional convolutions. The results are shown in Figure 10, where the source domain features after dimensionality reduction are represented as orange dots and the target domain features are represented as blue dots.
@@
The feature distribution of conventional convolution in the target domain is relatively scattered, and the overlap between the source and target domains is low, presenting a strong domain shift. The features learned by SAFDConvolution show a high degree of overlap between the source and target domain samples within the same cluster, indicating that this convolution effectively extracts domain-invariant and highly discriminative representations, reduces domain shift, and has strong generalization ability.

### v2

@@ Abstract
Deformable convolution can adaptively change the shape of convolution kernel by learning offsets to deal with complex shape features. We propose a novel plug-and-play deformable convolutional module that uses attention and feedforward networks to learn offsets, so that the deformable patterns can capture long-distance global features. Compared with previously existing deformable convolutions, the proposed module learns the sub-pixel displacement field and adaptively warps the feature maps across all channels (rather than directly deforms the convolution kernel), which is equivalent to a relative deformation of the kernel’s sampling grids, achieving global feature deformation and the decoupling of (kernel size)–(learning network). Considering that the fundus blood vessels have globally self-similar complex edges, we design a deep learning model for fundus blood vessel segmentation, GDCUnet, based on the proposed convolutional module. Empirical evaluations under the same configuration and unified framework show that GDCUnet has achieved state-of-the-art performance on public datasets. Further ablation experiments demonstrated that the proposed deformable convolutional module could more significantly learn the complex features of fundus blood vessels, enhancing the model’s representation and generalization capabilities. The proposed module is similar to the interface of conventional convolution, we suggest applying it to more machine vision tasks with complex global self-similar features. Our source code will be available at https://github.com/LexyZhu/GDCUnet.git
@@ 3.1 SAFDConvolution
[…]
@@
[…]
@@
The number of channels $C$ is usually divisible by 222, we denote that $M_{c}=C/2$ and split the last dimension:
@@
$\Delta\mathbf{p}(\mathbf{p}_{0}+\mathbf{p}_{n})$ $\xleftarrow{}\frac{1}{M_{c}}\sum_{m_{c}=1}^{M_{c}}\tilde{X}_{offset}$ $\in\mathbb{R}^{2}\text{ for each point in }B\times H\times W;$ (21)
@@ 3.2 GDCUnet
During the encoding stage, the Conv Block contains two $\times 3$ 3 conventional convolutions. SAFDConv Block uses two consecutive proposed SAFDConvolutions, where the hyperparameters can be set according to Table 1. Further, the Feature Incentive Block uses one $\times 7$ 7 conventional convolution.
@@
Figure 3: Overall pipeline of GDCUnet, where $C_{s}$ is 161616. Our network adopts a structure similar to Unet. In the encoding stage, Max-pooling is used to reduce the size of the feature map, and in the decoding stage, bilinear interpolation is used to restore the size of the feature map. We use tensor addition to achieve feature fusion of the feature maps from the previous stage. The modules using proposed SAFDConvolution are italicized bold.. To the best of our knowledge, the proposed deformable convolution is more suitable in the intermediate stage of feature extraction and decoding.
@@ 4.2 Implementation Details
Our experiments are built on CHASEDB1 [36], a public retinal blood vessel reference dataset. The dataset is divided into a training set and a test set in a ratio of 0.86:0.140.86:0.140.86 : 0.14. The experimental configuration summary is presented in Table 2. All experiments are conducted using the Pytorch framework [37], and the learning rate is uniformly optimized using cosine annealing. We evaluate the model on the test set after each epoch and report the highest score achieved.
@@
Input Resolution $\times 256$ 256
Output Resolution $\times 256$ 256
@@ 4.3 Metrics
Table 3: Summary of evaluation metrics, where IoU and Dice (F1 score) are the overall main indicators for evaluating the segmentation performance.
@@ 4.4 Main Results
Rolling-Unet #1.78M (#7.10M) [9] AAAI’2024 0.6008 (0.6082) 0.7506 (0.7563) 15.00 (16.00) 1.414 (1.732) 0.7745 (0.7549) 0.9828 (0.9826) 0.7568 (0.7578)
@@
Knowledge  [^2] We emphasize conducting experiments under the unified benchmark and configuration. However, the model based on prior knowledge characterizing the vascular topological structure involves separate settings for data processing, loss functions, etc., and experiments need to be conducted under separate configurations. DSCNet #- -M [34] ICCV’2023 0.6073 0.7521 15.00 2.000 0.7534 0.9832 0.7638
@@
Almost all GDCUNet Settings have achieved competitive performance. Setting 5 achieves the highest scores on the primary overall metrics IoU and Dice with only 1.40M parameters. At the same time, there is also more lightweight option (Setting 3) that achieves a balance between parameter quantity and segmentation performance. DconnNet has the highest recall but is limited overall. Benefit by the characterization ability of SAFDConvolution for globally self-similar complex local features, GDCUNet Setting 5 has reached state-of-the-art comprehensive level.
@@ 5.1 The Effect of SAFDConvolution
The main improvement made by GDCUNet is the addition of SAFDConvolution during the feature processing, thereby more effectively capturing complex local details with global self-similarity. Compared with other existing deformable convolution methods, SAFDConvolution can handle global features more effectively. We replace the convolution in the SAFDConvolution Block with other convolution (unify the kernel size $\times 5$ 5), and the quantitative evaluation results are shown in Table 5.
@@ 5.2 Feature Learning Significance
Figure 9: Distribution histograms of feature map grid values of conventional convolution. Please zoom in for a better illustration.
@@ 5.3 T-SNE Visualization
To further verify the generalization performance of SAFDConvolution, following the approach of J. Feng et al. [43], we use t-SNE to reduce the dimension of learned feature tensors. First, we visualize the training and testing features extracted by SAFDConv Block in GDCUnet. Then, we replace SAFDConvolutions with conventional convolutions while keeping other structures unchanged to visualize the training-testing features extracted by conventional convolutions. The results are shown in Figure 10, where the training features after dimensionality reduction are represented as orange dots and the testing features are represented as blue dots.
@@
The feature distribution of conventional convolution in the testing set is relatively scattered, and the overlap between the training and testing domains is low, presenting a strong domain shift. The features learned by SAFDConvolution show a high degree of overlap between the training and testing domain samples within the same cluster, indicating that this convolution effectively extracts domain-invariant and highly discriminative representations, reduces domain shift, and has strong generalization ability.